Sandy has 3.25 in change. She has five more dimes than quarters but twice as many nickels as dimes. Eecpress th given information as an algebraic exprssion and simplify.
Answer by Mathtut(3670)

let n,d,and q represent the number of nickels,dimes,and quarters,respectively
:
d=5q..................eq 1
n=2d..................eq 2
.05n+.1d+.25q=3.25....eq 3
:
re write eq 1 to q=d/5 and take that value and the value of n in eq 2 and plug them into eq 3
:
.05(2d)+.1d+.25(d/5)=3.25
:
.1d+.1d+.25(d/5)=3.25
:
.2d+.25(d/5)=3.25
:
d+.25d=16.25...multiplied all terms by 5 to get rid of fractional terms
:
1.25d=16.25
:
d=13
:
so now you can find out the values of n and q.